PC@LIVE wrote on 2024-06-09, 14:11:
I had an ASUS CUV4X in my hands, which needed to be tested, with a P3 370 733MHz, after the first attempts didn't go well, I understood that there was a RAM problem, I tried different SDRAMs, and the third module I tried, solved it the problem, the PC started, it's alive!!!!!!!!!!
While waiting to gather more hw, I tried booting from DOS, and launched a couple of programs, diagnostics and bench.
That heatsink seems like an odd one for a pentium III, is that maybe from some Socket 7 system?
I got hold of a Sony PRD-250 CD-ROM Discman player, which is a SCSI CD-ROM drive but is missing its cable and I've decided I like SCSI drives now. But it didn't read discs at all, just hunting around when trying to play pressed audio CDs.
Using a junky webcam, I confirmed that the laser was working and the focus system was working well, the carriage could move freely so it was probably a laser problem, like it's worn out.
I was getting nowhere and was going to give up then realised I have nothing to lose since they're not common drives and I had found the PRD-650 service manual which gave me some understanding of the TEST mode it could be put into by soldering together a solder jumper. I also learned that the laser assembly (DAX-02S) is probably irreplaceable.
I've never fixed a CD player before but I have two oscilloscopes, one is the Owon HDS2102S portable scope and the other is the Rigol DS1054Z - the latter gets used less because it is big and cumbersome in my opinion, but my view of it improved while trying to fix this drive. The control system is so much nicer than on a portable scope, and the screen is bigger.
So anyways, I took the player apart and started by taking pictures of everything - all the trimpots original positions and I tried marking them, this probably saved me.
The attachment IMG_3622 (Custom).JPG is no longer available
Because I hooked up the test points RFO (RF output?) to see the diamond pattern, VC as the ground point, TE for the E-F balance. Initially I tried using the Owon scope but I didn't know about its persistence settings so could not get anything resembling an 'eye' pattern that the service manual says to look for. Then settings were adjusted wildly for a while.
The attachment IMG_3624 (Custom).JPG is no longer available
I then switched to the Rigol scope, watched some videos of other people looking at this signal and started to understand what I was looking at. Soon after I could see something resembling the diamond type pattern but the voltage levels (vpp) were too low and it was noisy, so the laser did need boosting I think...?
Took off some capacitors to test whether those could be causing issues and no, they're not leaking and the ESR values were in spec for the 2 I tested.
Through the course of the night I adjusted everything, all the trimpots though only 3 mattered -
- The trimpot on the laser diode itself which of course the service manual didn't mention because that's set in the factory. It started off at something like 1.7v and I put it up to 1.71v or something - that was measuring on some test points by the laser diode so I'm pretty sure that's the voltage it's being fed?
- Then the Focus Bias adjustment which is the one that really made the difference, very finnicky adjustment
- The E-F Balance which seemed to set the high/low of the voltage and also made a difference to the eye pattern, in that the eye pattern only looked good once the E-F was set in the middle so the upper and lower waveforms matched and weren't clipping
There's also the Focus / Tracking gain adjustment pots but adjusting them did nothing positive so I pretty much have them back in their original positions which worked best.
I probably had the player functional last night since it was playing audio sometimes but broken up - that's going to happen in test mode because the carriage can be moved in/out using the |<< and >>| buttons, it's not playing an audio track, just a repeating bitstream.
Eventually I gave up thinking there was no hope and went to bed.
Got back to it in the morning and somehow I had the player reliably making sound in test mode, then playing audio and giving a nice 'diamond' / 'eye' pattern in something like 20 minutes of fiddling with it again.
The attachment prd-250-repair-eye-pattern.JPG is no longer available
From the above picture I realised the amplitude was now too high so put the laser voltage back down and the RFO test point now shows 1.2v vpp which is in spec I think. It's now adjusted just a little bit higher than its original setting.
The attachment prd-250-repair-eye-pattern-2.JPG is no longer available